3+3/4x=x/2 First you have to get ride of the denominator. You can do this by multiplying both sides of the equation by 4x: 4x(3+3/4x) = 4x(x/2)
Therefore 12x+3 =2x^2. This is a quadratic equation and you can rewrite it as
2x^2-12x-3= 0 using quadratic formula you get two answers one is x=(6+sqr root of 42)/2 and the other is x=(6-sqr root of 42)/2

Answer by Aldorozos(172)

3+3/4x=x/2
First you have to get ride of the denominator. You can do this by multiplying both sides of the equation by 4x: 4x(3+3/4x) = 4x(x/2)
Therefore 12x+3 =2x^2. This is a quadratic equation and you can rewrite it as
2x^2-12x-3= 0 using quadratic formula you get two answers one is x=(6+sqr root of 42)/2 and the other is x=(6-sqr root of 42)/2
See the quadratic equation below.